# Obfuscation Based Privacy Preserving Representations are Recoverable Using Neighborhood Information

### v1

@@ Abstract
Rapid growth in the popularity of AR/VR/MR applications and cloud-based visual localization systems has given rise to an increased focus on the privacy of user content in the localization process. This privacy concern has been further escalated by the ability of deep neural networks to recover detailed images of a scene from a sparse set of 3D or 2D points and their descriptors - the so-called inversion attacks. Research on privacy-preserving localization has therefore focused on preventing these inversion attacks on both the query image keypoints and the 3D points of the scene map. To this end, several geometry obfuscation techniques that lift points to higher-dimensional spaces, i.e., lines or planes, or that swap coordinates between points have been proposed. In this paper, we point to a common weakness of these obfuscations that allows to recover approximations of the original point positions under the assumption of known neighborhoods. We further show that these neighborhoods can be computed by learning to identify descriptors that co-occur in neighborhoods. Extensive experiments show that our approach for point recovery is practically applicable to all existing geometric obfuscation schemes. Our results show that these schemes should not be considered privacy-preserving, even though they are claimed to be privacy-preserving. Code will be available at https://github.com/kunalchelani/RecoverPointsNeighborhood.
@@ 1 Introduction
Visual localization estimates the position and orientation of a camera in a given scene and is central for autonomous navigation [66, 67], Simultaneous Localization and Mapping (SLAM) [14, 18], Augmented and Virtual Reality (AR/VR) [27, 50, 53], and Structure-from-Motion (SfM) [57, 58]. The best performing methods represent the scene with a 3D map, e.g., a Structure-from-Motion (SfM) point cloud [51, 54] or meshes [46]. To localize a given query image, they match the descriptors of 2D local features [40, 17] extracted from the query image against the descriptors of the 3D points in the map. The resulting 2D-3D point correspondences are used for camera pose estimation [23, 28, 33, 34, 35]. Such feature-based approaches are known to handle challenging conditions and to provide accurate pose estimates [68, 53]. However, they also pose a potential privacy risk because of inversion attacks [49]: it is possible to recover the query image in high detail from the 2D image features with inversion networks [49, 16]. One can also recover the map’s content from the 3D points and their descriptors [49, 62]. Thus, feature-based methods cannot be directly applied in settings where privacy is of concern [63, 64], such as when a user sends data to a localization service in the cloud or when 3D maps are stored on an external server.
Privacy-preserving localization methods aim to prevent such content recovery and mainly fall into two categories: descriptor obfuscation approaches define descriptors that are not informative enough to enable inversion but still allow accurate 2D-3D matching [22, 44, 48, 47]. Geometry obfuscation approaches replace each 2D or 3D point with a potentially infinite set of points [63, 64, 24, 25, 26, 36, 45, 60, 43]. For example, replacing a 3D point with a 3D line amounts to replacing the point with an infinite set of points forming the line passing through this point [63, 36] (see Fig. 1). By replacing points with infinite sets, these methods prevent a direct application of inversion attacks [49, 22]. In geometry obfuscation approaches, the function mapping a point to a set of points is carefully designed such that the resulting sets still enable pose estimation, e.g., the 2D-3D point matches are replaced by 2D point-to-3D line matches in [63].
@@
Figure 1: Geometry obfuscations allow the recovery of image details. The original point representations (row 1) are privacy revealing as full images can be recovered from them [49]. Hence, they are obfuscated using different schemes (row 2). In this paper (row 3), we show that given neighborhood information, it is possible to approximately recover the original point positions, again enabling image recovery (row 4).
@@
Geometry obfuscation methods are considered privacy-preserving by the community since it is unclear how to recover the original point positions from the obfuscations. However, none of the previous work proves that approximating the original point positions is impossible. On the contrary, [11] reveals the need for more scrutiny before claiming that a method is privacy-preserving: they show that when each 3D point is obfuscated by a line passing through the point with a random direction [63], it is possible to approximate the original 3D point position, thus enabling an inversion attack [49]. Their approach is based on two key insights: (1) the closest points on two such 3D lines are likely to be relatively close to the original 3D points. (2) to recover the position of a point $\mathbf{X}_{i}$ obfuscated by a line $\mathbf{l}_{i}$, it is important to know the neighbors of $\mathbf{l}_{i}$ (defined as the set of lines $\{\mathbf{l}_{j}\}$ corresponding to original 3D points that are nearest neighbors of $\mathbf{X}_{i}$). [11] cannot be applied to 2D point obfuscations [64, 45] and does not generalize since property (1) does not hold for all obfuscation schemes [64, 36, 45].
@@
Inspired by observation (2) from [11], we derive a novel and conceptually simple method for approximating the positions of the original points that is applicable for all of the currently proposed geometry obfuscation schemes [63, 64, 36, 45, 26, 43]. Our approach uses information about neighborhoods, i.e., about which obfuscated points correspond to nearest neighboring original points, and is computationally efficient [^1] In some cases, e.g.,  [36], our point recovery is faster to compute than obfuscating the points in the first place.. We show that the approximate point positions obtained with our method enable inversion attacks [49, 22]. Through extensive qualitative and quantitative experiments, we show that the proposed method is robust to errors in the neighborhoods, i.e., it is not necessary to have access to neighborhoods that contain only the nearest neighbors of the original points. We also demonstrate a simple way to learn neighborhoods from the descriptors associated with the geometric obfuscations, which need to be provided to enable visual localization. Our approach for computing neighborhoods serves as a proof-of-concept to show that our scheme for recovering points from geometric obfuscations is practically applicable.
@@
In summary, this paper makes the following contributions: (1) we present a novel framework for recovering approximate point positions from obfuscated scene representations. Our framework relies on neighborhood information and is applicable to all geometry obfuscation schemes from the literature. (2) we propose a learning-based approach for computing the required neighborhoods from the descriptors used for visual localization. (3) extensive experiments with neighborhoods provided by an oracle and our approach show the effectiveness of our recovery framework. Our results show that methods that are currently considered privacy-preserving do not in fact guarantee privacy. Our results highlight the need to derive clear conditions under which privacy can be guaranteed when proposing privacy-preserving localization approaches.
@@ 2 Related Work
Privacy Aspects of Visual Localization.  Cloud-based localization services require the exchange of information about the scene between the client and the server, leaving several cracks for possible privacy leaks. Naturally, such services should preserve the privacy of the 3D maps stored online from a curious/malicious server [16, 63, 45, 26, 11]. They should also preserve the client’s private information that is potentially sent to the server for localization, as part of the query image [64, 44, 16]. As noted in [26], even the server knowing the client’s accurate pose can potentially be a privacy risk. It is also shown that even the minimal requirements for running a robust localization service - returning the camera pose to the client - enable the approximate recovery of the scene layout by a malicious third party[12]. In this paper, we analyze to which extent privacy-preserving geometric obfuscations can reveal private content from a 3D map stored on a server and from a client’s query image.
@@ 4 Recovering Obfuscated Points using Neighborhood Information
Naturally, if the points $\mathcal{R}(\mathcal{O}(x_{j}))$ are close to the original points $x_{j}$ in the space $\mathbb{R}^{m}$, it can be expected that an inversion attack recovers a detailed image, potentially containing private information. Thus, if $d(\mathcal{R}(\mathcal{O}(x_{j})),x_{j})\leq\epsilon$, for some small $\epsilon$  [^4] Here, $d$ is the Euclidean distance in $\mathbb{R}^{m}$., the obfuscation $\mathcal{O}$ cannot be considered as privacy preserving. In this paper, we show that having information about the neighborhoods of the original points, we can compute a mapping $\mathcal{R}$ for which $\epsilon$ is sufficiently small for most points. Thus, private details can be identified in images recovered from the $\mathcal{R}(\mathcal{O}(x_{j}))$ and their descriptors.
Recovering points using neighborhood information.  Let us assume that for each obfuscated input point $\mathcal{O}(x_{j})$ we are given a set of neighbors $\mathcal{N}(\mathcal{O}(x_{j}))=\left\{\mathcal{O}(x_{j_{i}}):j_{i}\in[j_{1}\dots j_{K}]\right\}$. Here, $[j_{1}\dots j_{K}]$ are the indices of the K nearest neighbors (in $\mathbb{R}^{m}$) of the point $x_{j}$ among all points in ${P}$. In other words, the set $\mathcal{N}(\mathcal{O}(x_{j}))$ contains the obfuscated representations corresponding to the K nearest neighbors in $P$ of the original point $x_{j}$. Note that we do not assume that the original point positions are available. Furthermore, we assume that each original point $x_{j}$ is contained in its set $\mathcal{O}(x_{j})$, i.e., $\forall j$ $x_{j}\in\mathcal{O}(x_{j})$. This assumption holds for approaches that map $x$ to lines [63, 64, 36] or planes [26] passing through $x$. It does not hold for coordinate permutation-based obfuscation. However, as detailed below, for this case we can extend $\mathcal{O}(x_{j})$ to include all points on $m$ lines passing through $\mathcal{O}(x_{j})$ as one of them contains $x_{j}$.
@@
Here, $d(\mathcal{O}(x_{j_{i}}),x)$ is the Euclidean distance of a point $x\in\mathbb{R}^{m}$ from its closest point in $\mathcal{O}(x_{j_{i}})$. Note that the point $\mathcal{R}(\mathcal{O}(x_{j}))$ that minimizes (4) can be far away from the true point position $x_{j}$. However, in our experience, using sufficiently many neighbors "pulls" $\mathcal{R}(\mathcal{O}(x_{j}))$ towards $x_{j}$. Also note that we solve (4) per point $x_{j}$, rather than taking point estimates for the neighbors into account. This makes our recovery approach parallelizable.
@@ 5 Estimating Neighborhoods From Descriptors
Computing the recovered points $\mathcal{R}(\mathcal{O}(x_{j}))$ from the obfuscations $\mathcal{O}(x_{j})$ using (4) assumes that we have information about the neighbors of each original point $x_{j}$. In  [11], such a neighborhood is geometrically estimated by using the distance between pairs of 3D lines as a proxy for the distance between the original points. However, their approach requires that the line directions are random and that lines are thus unlikely to intersect in 3D. This assumption does not hold for 2D lines and orthogonal 3D planes.
In the context of visual localization, each obfuscated point is associated with a descriptor that is used for feature matching between the 2D image query and the 3D map. We use these descriptors to estimate the required neighborhoods.
Intuitively, local structures (captured by the K nearest neighbors of a point) are not unique for each scene, but similar-looking structures can be found in other scenes. This motivates our learning-based approach for estimating the neighborhoods. Given enough scenes as training data, we let a neural network learn about such patterns, which in turn can be used to determine neighborhoods. We pose the task of neighborhood estimation as a feature matching task [52, 39, 65]: given a set of descriptors, we learn a similarity score between all pairs of descriptors that is inversely proportional to the distance between the original points. More specifically, the network takes as input the descriptors and outputs a row-normalized similarity matrix with high entries between the points that are likely to correspond to neighboring points. The network is made of several self-attention blocks [70] that draw contextual cues between the descriptors. It is trained in a supervised manner with the binary cross-entropy loss. The entry $(i,j)$ of the similarity matrix forms a positive if the $j^{\text{th}}$ point is within the $K$ closest points to the $i^{\text{th}}$ point.
@@
7-scenes Cambridge 7-scenes Cambridge
@@
Table 1: Geometric accuracy of recovery from obfuscation of 2D SIFT [40] points. Percentage of points recovered within error thresholds using simulated neighborhoods with different inlier ratios. Images sizes 7-scenes [61] - $640\times 480$ and Cambridge[31] - $1024\times 576$ pixels.
The network training is not tied to a dataset or to a specific point distribution but only to the type of local descriptor. Thus, the network can be trained on any data where pairs of points and descriptors are available, and it can then be applied to the descriptors of the obfuscated representations. As shown in the experiments, such a simple network predicts neighborhoods that are sufficiently reliable to allow the proposed recovery method to reveal private content robustly. Note that our approach is intended as a proof-of-concept to show that our attack from Sec. 4 is practically feasible. We believe that better results can be obtained by tuning the network architecture and using larger training sets. However, such optimizations are outside the scope of this work.
6 Experimental Evaluation
@@
7-scenes Cambridge 7-scenes Cambridge 7-scenes Cambridge 7-scenes Cambridge
@@
Table 2: Geometric accuracy of recovery from obfuscation of 3D points triangulated from SIFT [40] features. Ratio of points recovered within error thresholds from oracle neighborhoods with various inlier ratios (In.) on the indoor 7-scenes [61] and the outdoor Cambridge [31] datasets. The line obfuscations PPL [36] and [43] are more susceptible to point recovery as compared to plane [26] and point-permutation [45] obfuscations.
We evaluate the recovery method based on how well it recovers the points obfuscated by 6 different obfuscation schemes in 3D - random lines [63] (OLC), two variants of paired points lines [36] (PPL and PPL+), ray clouds [43], planes [26], and Coordinate Permutation (CP) [45]. In 2D, we evaluate the recovery from points lifted to random lines [64] and with CP [45]. We experiment with two types of descriptors: the hand-crafted SIFT [40] and the learning-based SuperPoint [17]. To visually assess the information revealed by the point recovery, we further invert the recovered points and their descriptors into images of the scene with an inversion network [49, 22]. For most of our analysis, we use oracle-provided neighborhoods, i.e., neighborhoods directly obtained from the nearest neighbors of the original points, rather than neighborhoods computed using our approach from Sec. 5. This provides us full control over the quality of assumed neighborhoods, i.e., the inlier ratios, which is well suited for our analysis. This also allows us to showcase the robustness of our approach. Finally, to show the practical feasibility of the attack, we present results with neighborhoods estimated by our learning-based approach. The point recovery on these estimated neighborhoods reveals private content even though the proposed neighborhood network is only a proof of concept. As such, it is simple and has limited scalability to a few thousand descriptors whereas point clouds usually involve several hundred thousand descriptors. Therefore we run this end-to-end evaluation only in 2D. This network however shows the potential that a similar architecture could be trained for more descriptors given hardware with enough memory.

### v2

@@ Abstract
The rapid growth of AR/VR/MR applications and cloud-based visual localization has heightened concerns over user privacy. This privacy concern has been further escalated by the ability of deep neural networks to recover detailed images of a scene from a sparse set of 3D or 2D points and their descriptors - the so-called inversion attacks. Research on privacy-preserving localization has therefore focused on preventing such attacks through geometry obfuscation techniques like lifting points to higher dimensions or swapping coordinates. In this paper, we reveal a common vulnerability in these methods that allows approximate point recovery using known neighborhoods. We further show that these neighborhoods can be computed by learning to identify descriptors that co-occur in neighborhoods. Extensive experiments demonstrate that all existing geometric obfuscation schemes remain susceptible to such recovery, challenging their claims of being privacy-preserving. Code will be available at https://github.com/kunalchelani/RecoverPointsNeighborhood.
@@ 1 Introduction
Visual localization estimates the position and orientation of a camera in a given scene and is central for autonomous navigation [65, 66], Simultaneous Localization and Mapping (SLAM) [13, 17], Augmented and Virtual Reality (AR/VR) [26, 49, 52], and Structure-from-Motion (SfM) [56, 57]. The best performing methods represent the scene with a 3D map, e.g., a Structure-from-Motion (SfM) point cloud [50, 53]. To localize a given query image, they match the descriptors of 2D local features [39, 16] extracted from the query image against the descriptors of the 3D points in the map. The resulting 2D-3D point correspondences are used for camera pose estimation [22, 27, 32, 33, 34]. Such feature-based approaches are known to handle challenging conditions and to provide accurate pose estimates [67, 52]. However, they also pose a potential privacy risk because of inversion attacks [48]: it is possible to recover the query image in high detail from the 2D image features with inversion networks [48, 15]. One can also recover the map’s content from the 3D points and their descriptors [48, 61]. Thus, feature-based methods cannot be directly applied in settings where privacy is of concern [62, 63], such as when a user sends data to a localization service in the cloud or when 3D maps are stored on an external server.
@@
Figure 1: Geometry obfuscations allow the recovery of image details. The original point representations are privacy revealing as full images can be recovered from them [48]. Different obfuscation schemes are used to modify them. In this paper, we show that given neighborhood information, it is possible to approximately recover the original point positions, again enabling image recovery.
Privacy-preserving localization methods aim to prevent content recovery and mainly fall into two categories: descriptor obfuscation approaches, which modify descriptors to prevent inversion while enabling accurate 2D-3D matching [21, 43, 47, 46], and geometry obfuscation approaches, which replace each 2D or 3D point with a potentially infinite set of points [62, 63, 23, 24, 25, 35, 44, 59, 42]. An example of geometry obfuscation is lifting points to lines, which replaces each point with an infinite set of points lying on a line through the corresponding point [62, 35] (see Fig. 1). By substituting points with potentially infinite sets, these methods prevent the direct application of inversion attacks [48, 21]. Geometric obfuscation approaches carefully design the function mapping a point to a set of points so that the resulting sets still enable pose estimation; for instance, in [62], 2D-3D point matches are replaced by 2D point-to-3D line matches.
@@
Geometry obfuscation methods are considered privacy-preserving by the community since it is unclear how to recover the original point positions from the obfuscations. However, none of the previous work proves that approximating the original point positions is impossible. On the contrary, [10] reveals the need for more scrutiny before claiming that a method is privacy-preserving: they show that when each 3D point is obfuscated by a line passing through the point with a random direction [62], it is possible to approximate the original 3D point position, thus enabling an inversion attack [48]. Their approach is based on two key insights: (1) the closest points on two such 3D lines are likely to be relatively close to the original 3D points. (2) to recover the position of a point $\mathbf{X}_{i}$ obfuscated by a line $\mathbf{l}_{i}$, it is important to know the neighbors of $\mathbf{l}_{i}$ (defined as the set of lines $\{\mathbf{l}_{j}\}$ corresponding to original 3D points that are neighbors of $\mathbf{X}_{i}$). [10] cannot be applied to 2D point obfuscations [63, 44] and does not generalize since property (1) does not hold for all obfuscation schemes [63, 25, 44].
@@
Inspired by observation (2) from [10], we derive a novel and conceptually simple method for approximating the positions of the original points that is applicable for all of the currently proposed geometry obfuscation schemes [62, 63, 35, 44, 25, 42]. Our approach uses information about neighborhoods, i.e., about which obfuscated points correspond to nearest neighboring original points, and is computationally efficient [^1] In some cases, e.g.,  [35], our point recovery is faster to compute than obfuscating the points in the first place.. We show that the approximate point positions obtained with our method enable inversion attacks [48, 21]. Through extensive qualitative and quantitative experiments, we demonstrate that the proposed method is robust to errors in neighborhoods, i.e., it does not require access to neighborhoods containing only the nearest neighbors of the original points. Additionally, we present a simple approach for learning neighborhoods from the descriptors associated with geometric obfuscations, which must be provided to enable visual localization. Our method for computing neighborhoods serves as a proof of concept, demonstrating that our scheme for recovering points from geometric obfuscations is practically applicable.
@@
In summary, this paper makes the following contributions: (1) we present a novel framework for recovering approximate point positions from obfuscated scene representations. Our framework relies on neighborhood information and is applicable to all geometry obfuscation schemes from the literature. (2) we propose a learning-based approach for computing the required neighborhoods from the descriptors used for visual localization. (3) extensive experiments with neighborhoods provided by an oracle and our approach show the effectiveness of our recovery framework. Our results show that methods that are currently considered privacy-preserving do not in fact guarantee privacy, and highlight the need to derive clear conditions under which privacy can be guaranteed when proposing privacy-preserving localization approaches.
@@ 2 Related Work
Privacy Aspects of Visual Localization.  Cloud-based localization services require the exchange of information about the scene between the client and the server, leaving several cracks for possible privacy leaks. Naturally, such services should preserve the privacy of the 3D maps stored online from a curious/malicious server [15, 62, 44, 25, 10]. They should also preserve the client’s private information that is potentially sent to the server for localization, as part of the query image [63, 43, 15]. As noted in [25], even the server knowing the client’s accurate pose can potentially be a privacy risk. It is also shown that even the minimal requirements for running a robust localization service - returning the camera pose to the client - enable the approximate recovery of the scene layout by a malicious third party[11]. In this paper, we analyze the extent to which privacy-preserving geometric obfuscations can reveal private content from a 3D map stored on a server and from a client’s query image.
@@ 4 Recovering Obfuscated Points using Neighborhood Information
Naturally, if the points $\mathcal{R}(\mathcal{O}(x_{j}))$ are close to the original points $x_{j}$ in the space $\mathbb{R}^{m}$, it can be expected that an inversion attack recovers a detailed image, potentially containing private information. Thus, if $d(\mathcal{R}(\mathcal{O}(x_{j})),x_{j})\leq\epsilon$, for some small $\epsilon$  [^4] Here, $d$ is the Euclidean distance in $\mathbb{R}^{m}$., the obfuscation $\mathcal{O}$ cannot be considered as privacy preserving. In this paper, we show that having information about the neighborhoods of the original points, we can compute a mapping $\mathcal{R}$ for which $\epsilon$ is sufficiently small for most points. Thus, private details can be identified in images recovered from the $\mathcal{R}(\mathcal{O}(P))$.
Recovering points using neighborhood information.  Let us assume that for each obfuscated input point $\mathcal{O}(x_{j})$ we are given a set of neighbors $\mathcal{N}(\mathcal{O}(x_{j}))=\left\{\mathcal{O}(x_{j_{i}}):j_{i}\in[j_{1}\dots j_{K}]\right\}$. Here, $[j_{1}\dots j_{K}]$ are the indices of the K nearest neighbors (in $\mathbb{R}^{m}$) of the point $x_{j}$ among all points in ${P}$. In other words, the set $\mathcal{N}(\mathcal{O}(x_{j}))$ contains the obfuscated representations corresponding to the K nearest neighbors in $P$ of the original point $x_{j}$. The assumption that each original point $x_{j}$ is contained in its set $\mathcal{O}(x_{j})$, i.e., $\forall j$ $x_{j}\in\mathcal{O}(x_{j})$ holds for approaches that map $x$ to lines [62, 63, 35, 42] or planes [25] passing through $x$. It does not hold for coordinate permutation-based obfuscation. However, as detailed below, for this case we can extend $\mathcal{O}(x_{j})$ to include all points on $m$ lines passing through $\mathcal{O}(x_{j})$ as one of them contains $x_{j}$.
@@
Here, $d(\mathcal{O}(x_{j_{i}}),x)$ is the Euclidean distance of a point $x\in\mathbb{R}^{m}$ from its closest point in $\mathcal{O}(x_{j_{i}})$. Note that the point $\mathcal{R}(\mathcal{O}(x_{j}))$ that minimizes (4) can be far away from the true point position $x_{j}$. However, in our experience, using sufficiently many neighbors ”pulls” $\mathcal{R}(\mathcal{O}(x_{j}))$ towards $x_{j}$. Also note that we solve (4) per point $x_{j}$, rather than taking point estimates for the neighbors into account. This makes our recovery approach parallelizable.
@@ 5 Estimating Neighborhoods From Descriptors
Computing the recovered points $\mathcal{R}(\mathcal{O}(x_{j}))$ from the obfuscations $\mathcal{O}(x_{j})$ using (4) assumes that we have information about the neighbors of each original point $x_{j}$. In  [10], such a neighborhood is geometrically estimated by using the distance between pairs of 3D lines as a proxy for the distance between the original points. However, their approach requires that the line directions are random and that lines are thus unlikely to intersect in 3D. This assumption does not hold for 2D lines, orthogonal 3D planes and ray clouds [42].
In the context of visual localization, each obfuscated point is associated with a descriptor that is used for matching the 2D image query with the 3D map points. We use these descriptors to estimate the required neighborhoods.
Intuitively, local structures (captured by the neighbors of a point) are not unique for each scene, but similar-looking structures can be found in other scenes. This motivates our learning-based approach for estimating the neighborhoods. Given enough scenes as training data, we let a neural network learn about such patterns, which in turn can be used to determine neighborhoods. We pose the task of neighborhood estimation as a feature matching task [51, 38, 64]: given a set of descriptors, we learn a similarity score between all pairs of descriptors that is inversely proportional to the distance between the original points. More specifically, the network takes as input the descriptors and outputs a row-normalized similarity matrix with high entries between the points that are likely to correspond to neighboring points. The network is made of several self-attention blocks [69] that draw contextual cues between the descriptors. It is trained in a supervised manner with the binary cross-entropy loss. The entry $(i,j)$ of the similarity matrix is positive if the $j^{\text{th}}$ point is within the $K$ closest points to the $i^{\text{th}}$ point.
@@
7-scenes [60] Cambridge [30] 7-scenes [60] Cambridge [30]
@@
Table 1: Geometric accuracy of recovery from obfuscation of 2D SIFT [39] points. Percentage of points recovered within error thresholds using oracle neighborhoods with different inlier ratios. Sizes: 7-scenes [60] - $640\times 480$ and Cambridge [30] - $1024\times 576$.
@@
7-scenes [60] Cambridge [30] 7-scenes [60] Cambridge [30] 7-scenes [60] Cambridge [30] 7-scenes [60] Cambridge [30]
@@
Table 2: Geometric accuracy of recovery from obfuscation of 3D points (suing SIFT [39]). Ratio of points recovered within error thresholds from oracle neighborhoods with different inlier ratios (In.) on 7-scenes [60] and Cambridge [30] datasets. The line obfuscations PPL [35] and [42] are more susceptible to point recovery as compared to plane [25] and point-permutation [44] obfuscations.
The network training is only tied to the descriptor, and thus, a network can be trained on any data where point positions and associated descriptors are available, As shown in the experiments, such a simple network predicts neighborhoods that are sufficiently reliable to allow the proposed recovery method to reveal private content robustly. Note that our approach is intended as a proof-of-concept to show that our attack from Sec. 4 is practically feasible. We believe that better results can be obtained by tuning the network architecture and using larger training sets. However, such optimizations are outside the scope of this work.
6 Experimental Evaluation
We evaluate the recovery method based on how well it recovers the points obfuscated by 6 different obfuscation schemes in 3D - random lines [62] (OLC), two variants of paired points lines [35] (PPL and PPL+), the default ray clouds [42], planes [25], and Coordinate Permutation (CP) [44]. In 2D, we evaluate the recovery from points obfuscated using random lines [63] and with CP [44]. We experiment with two widely used descriptors: the hand-crafted SIFT [39] and the learning-based SuperPoint [16]. To visually assess the information revealed by the point recovery, we further invert the recovered points and their descriptors into images of the scene with an inversion network [48, 21]. For most of our analysis, we use oracle-provided neighborhoods, i.e., neighborhoods directly obtained from the nearest neighbors of the original points, rather than neighborhoods computed using our approach from Sec. 5. This provides us full control over the quality of assumed neighborhoods, i.e., the inlier ratios, which is well suited for our analysis. This also allows us to showcase the robustness of our approach. Finally, to show the practical feasibility of the attack, we present results with neighborhoods estimated by our learning-based approach. The point recovery on these estimated neighborhoods reveals private content even though the proposed neighborhood network is only a proof of concept. As such, it is simple and has limited scalability to a few thousand descriptors whereas point clouds usually involve several hundred thousand descriptors. Therefore we run this end-to-end evaluation only in 2D. This network however shows the potential that a similar architecture could be trained for more descriptors given hardware with enough memory.